Please answer both parts
A continuous stream contains 40 mol % benzene and
Answer #1

1. Benzene= 40%

this indicates mass of benene = 40 X 78 (where 78 is mass of benzene)

= 3120g

Toulene = 60%

this indicates mass of toulene = 60 X 92 (where 92 is mass of toulene)

= 5520g

massfraction of benzene = mass of benene/ total mass

= 3120/(3120+5520) = 0.361

massfraction of toulene = mass of toulene/ total mass

= 5520/(3120+5520) = 0.639